OpenStudy (anonymous):

Fruit wingspantail comes in large and small cylindrical cans. The larger can has a radius and height that are both twice as long as that of the radius and height of the smaller can. If the volume of the smaller can is 28.64 in3, what is the volume of the larger can? A. 171.84 in3 B. 114.56 in3 C. 57.28 in3 D. 229.12 in3

OpenStudy (anonymous):

\[Equation of a cylinder : \pi * r^2 * h\]

OpenStudy (anonymous):

So the smaller can has a volume of 28.64 in^3 We can assume that \[\pi * r^2 * h = 28.64 \]

OpenStudy (anonymous):

We know that the larger can has twice the radius and twice the height. r=2r Volume of large sphere=\[\pi* (2r)^2 * 2h\] \[\pi* 4r*2h = ??\]

OpenStudy (anonymous):

\[\pi* r^2 * h =28.64\] \[\pi* 4r^2 *2h =????\]

OpenStudy (anonymous):

is the answer D

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what is the difference between the small cans volume and the larger one?

OpenStudy (anonymous):

π∗r2∗h=28.64 π∗4r2∗2h=????